-8x - 5y = -45 as a function of x.
Options:
A)f(x)=8x + 9
B)f(x) = -8/5 + 9
c)f(x) = 5/8x - 45
d)f(x) = -8x + 45

Found 2 solutions by fractalier, tutorcecilia:
Answer by fractalier(6550)

To find y as a function of x, we merely solve for y in terms of x...thus from
-8x - 5y = -45
we add 8x to both sides, then divide both sides by -5...
-5y = 8x - 45
y = (-8/5)x + 9
I imagine this to be Choice B but you typed it wrong (I edited it)

Answer by tutorcecilia(2152)


A function of x can be written as f(x) or y. So you are solving for y.
-8x - 5y = -45
-8x+8x-5y=+8x-45
-5y=+8x-45
.
-5y/-5=+8x/-5-45/-5
.
y=(-8/5)x+9